# Butter: Frequency Consistency and Hierarchical Fusion for Autonomous Driving Object Detection

### v1

@@ 2. Related Works
Figure 2. Comprehensive workflow of the Butter model for autonomous driving object detection. (1) The workflow begins with a monocular image of $640\times 640$ pixels processed through the HGStem of the Backbone to extract features. These features are subsequently refined through a series of lightweight HGBlocks, Depthwise Convolutions (DWConv) (Howard et al., 2017), and Convolutional Block Attention Modules (CBAM) (Woo et al., 2018) before entering the Neck module. The Neck consists of two parts: FAFCENet and PHFFNet. Following the Neck, the model utilizes four heads in the Head layer to produce outputs, including the original image annotated with class labels, confidence scores, and bounding boxes. (2) The CBAM module in the lower left corner applies channel and spatial attention to guide the network toward informative features. (3) The Hierarchical Fusion Block in the upper right corner enables multi-level feature interaction within the Context-Aware Spatial Fusion (CASF) module. Horizontal arrows represent feature exchange, while only diagonal arrows are used to indicate upsampling and downsampling.
@@ 3.1. Overview
Overall Framework. As shown in Fig.2, Butter consists of three primary branches. The Backbone Branch (Fig.2 (a)) leverages a lightweight version of the HGNetV2 architecture, enhanced with lightweight Depthwise Convolutions (DWConv) (Howard et al., 2017) and Convolutional Block Attention Modules (CBAM) (Woo et al., 2018) Module. This branch processes input monocular $640\times 640$ RGB images to extract essential features, improving both computational efficiency and detection accuracy. The Neck Branch (Fig. 2 (b)) consists of two crucial elements: FAFCE as a component and PHFFNet as a module. The FAFCE component enhances the model’s ability to maintain feature consistency and boundary precision. At the same time, the PHFFNet module integrates low-level and high-level features progressively, helping to bridge semantic gaps and improve the accuracy of object detection. The Head Branch (Fig. 2 (c)) distinguishes itself from traditional YOLOv12 (Tian et al., 2025) models by using four detection heads instead of the typical three.
@@ 3.3. Frequency-Adaptive Feature Consistency Enhancement (FAFCE) Component
where $\mathbf{A}^{l}\in\mathbb{R}^{2H\times 2W\times C}$ represents the $l$-th feature map produced by the backbone, and $\mathbf{B}^{l+1}\in\mathbb{R}^{H\times W\times C}$ denotes the fused feature at the $(l+1)$-th level. The operator $\mathcal{U}_{\theta}^{\mathrm{UP}}$ refers to an upsampling function and $\theta$ is the learnable parameter. The two feature maps are assumed to have the same number of channels; otherwise, a $1\times 1$ convolution can be applied as a projection function, which we omit here for simplicity. Despite its simplicity, this fusion approach causes semantic inconsistency and boundary misalignment. These issues harm dense prediction in autonomous driving. Simple interpolation spreads errors spatially, which leads to misclassifications. It also oversmooths outputs, weakening boundary localization. Additionally, the approach underutilizes boundary cues from lower-level features, which are vital for accurate detection in complex environments.
@@
where $Re_{(u,v)}$ is the resampling operation with displacement $(u,v)$, used to re-align the features from the previous stage, $\mathbf{B}_{Initial}^{l+1}$ is the refined feature from the stage 1, and $\widetilde{\mathbf{B}}^{l+1}$ represents the feature map from the amplifier. Stage 2 refines the spatial alignment of the features, ensuring that they are better aligned for the final fusion stage.
@@ 3. Final Feature Fusion
Let $x_{n}^{\rightarrow l,ij}$ represents the feature vector at spatial position $(i,j)$ from level $n$ to level $l$, and let $y_{l}^{ij}$ denotes the resulting fused feature vector at position $(i,j)$ and level $l$. We integrate feature vectors from 3 distinct Hierarchical Fusion Block as illustrated in Fig. 2. The fusion is performed by summing over these levels, with $\lambda_{n}^{ij}$ representing the dynamic spatial weight for the feature vector at level $n$ and spatial position $(i,j)$. These weights dynamically adjust based on the significance of the features at each spatial location, allowing the fusion process to emphasize relevant features and suppress less important ones.

### v2

@@ 2. Related Works
Figure 2. Comprehensive workflow of the Butter model for autonomous driving object detection. (1) The workflow begins with a monocular image of $\times 640$ 640 pixels processed through the HGStem of the Backbone to extract features. These features are subsequently refined through a series of lightweight HGBlocks, Depthwise Convolutions (DWConv) (Howard et al., 2017), and Convolutional Block Attention Modules (CBAM) (Woo et al., 2018) before entering the Neck module. The Neck consists of two parts: FAFCENet and PHFFNet. Following the Neck, the model utilizes four heads in the Head layer to produce outputs, including the original image annotated with class labels, confidence scores, and bounding boxes. (2) The CBAM module in the lower left corner applies channel and spatial attention to guide the network toward informative features. (3) The Hierarchical Fusion Block in the upper right corner enables multi-level feature interaction within the Context-Aware Spatial Fusion (CASF) module. Horizontal arrows represent feature exchange, while only diagonal arrows are used to indicate upsampling and downsampling.
@@ 3.1. Overview
Overall Framework. As shown in Fig.2, Butter consists of three primary branches. The Backbone Branch (Fig.2 (a)) leverages a lightweight version of the HGNetV2 architecture, enhanced with lightweight Depthwise Convolutions (DWConv) (Howard et al., 2017) and Convolutional Block Attention Modules (CBAM) (Woo et al., 2018) Module. This branch processes input monocular $\times 640$ 640 RGB images to extract essential features, improving both computational efficiency and detection accuracy. The Neck Branch (Fig. 2 (b)) consists of two crucial elements: FAFCE as a component and PHFFNet as a module. The FAFCE component enhances the model’s ability to maintain feature consistency and boundary precision. At the same time, the PHFFNet module integrates low-level and high-level features progressively, helping to bridge semantic gaps and improve the accuracy of object detection. The Head Branch (Fig. 2 (c)) distinguishes itself from traditional YOLOv12 (Tian et al., 2025) models by using four detection heads instead of the typical three.
@@ 3.3. Frequency-Adaptive Feature Consistency Enhancement (FAFCE) Component
[…]
@@
[…]
@@ 3. Final Feature Fusion
[…]